OpenStudy (anonymous):

2(x + 6) − (x + 1) = 16

OpenStudy (anonymous):

its 5 for sure

OpenStudy (anonymous):

i think

OpenStudy (mayankdevnani):

2x+12-x-1=16 x-11=16 x=16+11=27

OpenStudy (mayankdevnani):

ok @rhondaycole

OpenStudy (anonymous):

so are you saying im wrong?

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

@mayankdevnani is r8...

OpenStudy (anonymous):

2(x+6) - (x+1) = 16 2x+12-x-1=16 x+11=16 x=16-11 x=5

OpenStudy (anonymous):

i say your wrong

OpenStudy (anonymous):

thank you @LMar100%

OpenStudy (anonymous):

im right

OpenStudy (anonymous):

hey @LMar100% ur 3rd step is wrong... it is x-11

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

so its 5?

OpenStudy (mayankdevnani):

sry.. m wrong

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

thnx

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im right yay

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

no its not 5... if it is so pls tel me how

OpenStudy (anonymous):

it is 5 i know

OpenStudy (anonymous):

can u explain d procedure... @Bgarza95

OpenStudy (anonymous):

no i can not try it on mathway.com that will explain it to you

OpenStudy (anonymous):

2(x+6) - (x+1) = 16 expand the brackets 2x+12-x-1=16 now if you look the -1 is multiplied by each term in the second bracket x+11=16 then you subtract 11 from both sides x=16-11 x=5

OpenStudy (anonymous):

thanks @LMar100%

OpenStudy (anonymous):

no problem